Circular No. 8378

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dj IN NGC 2403 F. Patat, European Southern Observatory; S. Benetti and A. Pas-

torello, Istituto Nazionale di Astrofisica; and A. V. Filippenko, University of California, Berkeley, on behalf of the ERTN (see IAUC 8341), report that a fully reduced spectrum (range 320–900 nm; resolution 1.2 nm) of SN 2004dj (cf. IAUC 8377), obtained on Aug. 3.17 UT by J. Aceituno at the Calar Alto Observatory 2.2-m telescope (+ CAFOS), shows it to be a normal type-II-P supernova, very similar to SN 1999em, ∼ 3 weeks after explosion (Leonard et al. 2001, Ap.J. 553, 861; Hamuy et al. 2001, Ap.J. 558, 615). Besides a very-well-developed hydrogen Balmer series, lines from Na I, Fe II, and Ca II are clearly visible, all with P-Cyg profiles. The Na I D feature is possibly blended with He I 587.6-nm. The spectrum also shows an unresolved absorption at 589.7 nm, identified as interstellar Na I D, most likely generated by intervening material in the host galaxy. The equivalent width of this feature, 0.11 nm, corresponds to E(B − V) ∼ 0.18 mag. The photospheric expansion velocities, deduced from the minima of absorption troughs of various lines, are as follows: Hα, 6700 km/s; Hβ, 5500; Fe II 516.9-nm, 4150; Fe II 501.8-nm, 3800; Fe II 492.4-nm, 3850 km/s.

COMETS C/2001 Q4 (NEAT) AND C/2003 K4 (LINEAR) C. E. Woodward and M. S. Kelley, University of Minnesota; and D. H.

Wooden, Ames Research Center, NASA, report 8–13-μm spectrophotometry of these comets using the NASA Ames HIFOGS spectrometer at the Infrared Telescope Facility 3-m reflector: “Weak silicate-feature emission (cf. IAUC 8360, 8339) is present in the 10-μm spectra of C/2001 Q4 on July 28.24 UT, when the observed N-band magnitude (3′′ circular aperture) was 3.7 ± 0.4. Preliminary analysis of the featureless 10-μm spectra of C/2003 K4 suggests that large amorphous carbon and silicate grains (radius > ∼ 0.7 μm) dominate the coma. Further to IAUC 8361, no structure attributable to crystalline silicates was evident. The observed N-band magnitudes (3′′ circular aperture) of C/2003 K4 were: July 26.24, 3.6 ± 0.4; 27.24, 3.4 ± 0.2.”

Visual total-magnitude estimates of C/2003 K4: June 27.11 UT, 7.6 (J. J. Gonzalez, Leon, Spain, 7×50 binoculars); July 12.94, 6.8 (A. Baransky, Bucha, Ukraine, 10×50 binoculars); 16.96, 7.1 (A. Kammerer, Malsch, Germany, 9×63 binoculars); 20.49, 6.8 (Y. Nagai, Nagano, Japan, 7×35 binoculars); 28.93, 6.7 (A. Diepvens, Balen, Belgium, 20×50 binoculars); Aug. 1.92, 6.5 (J. G. de S. Aguiar, Campinas, Brazil, 11×80 binoculars).

2004 August 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green